OpenStudy (anonymous):

All of the following are equivalent, except _____. 2x + x x(2 + 1) 2x² 3x

OpenStudy (anonymous):

\[2x ^{2}\]

OpenStudy (simplymarie_x):

^

OpenStudy (faiqraees):

@livingmaybe Please resist from giving direct answers. It is against the code of conduct and does not help anyone learn. Plus it discourages other inspirational teachers to teach properly since the provision of the answer before the solution, makes the questioner to skip the solution part. It will be very helpful to the questioner and even the openstudy community that you try to walk them through the solution and explain them. For reference visit http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

you were wrong @FaiqRaees

OpenStudy (anonymous):

I already got he answer thanks :)

OpenStudy (mathstudent55):

What answer did you get?

OpenStudy (anonymous):

Uh no you weren't its @2x^2

OpenStudy (anonymous):

all of them have variable x

OpenStudy (mathstudent55):

2x + x = 2x + 1x = 3x x(2 + 1) = x* 3 = 3x 2x² = 2x² 3x = 3x

OpenStudy (mathstudent55):

Three expressions equal 3x, and the other one equals 2x^2.

OpenStudy (anonymous):

yeah exactly now get off my question please byeee

OpenStudy (anonymous):

if someone closes their question then they got the answer

OpenStudy (anonymous):

thanks